### v1

@@ Appendix A Additional Experiment Results
Appendix B Introduction to Adversarial Samples, Adversarial Training and Mixup
B.1 Adversarial Samples and Adversarial Training
Adversarial sample are used in evaison attacks to deceive neural networks. They can be generated by taking small steps in the direction of the loss gradient Goodfellow et al. (2014). A commonly used methods to generate adversarial samples is Projected Gradient Descent Attack (PGD) Kurakin et al. (2016). PGD searches for adversarial samples interactively in a small area around the input by adding perturbations following loss gradients:
$x^{adv}_{t+1}=x^{adv}_{t}+\alpha\cdot sign(\nabla_{x}\mathcal{L}(\mathbb{F}(x^{adv}_{t}),y))$ (11)
Adversarial training (AT) was first proposed by PGD-AT to increase the model’s robustness to adversarial samples Madry et al. (2018). Adversarial robustness prevents neural networks from being fooled, and can be achieved by ensuring that all samples near the sample $x$ have the same classification with $x$: $\mathbb{F}(x)=\mathbb{F}(x+\delta)$ for $\forall|\delta|<\epsilon$. AT introduces adversarial samples to the training set, seeking the worst case in scope $\ B_{\epsilon}(x)=\{x^{\prime}:c_{\mathcal{X}}(x,x^{\prime})\leq\epsilon)\}$:
$\inf_{\theta}\mathbb{E}_{\mathbb{P}}[\sup_{x^{\prime}\in B_{\epsilon}(x)}\mathcal{L}(\mathbb{F}(x^{\prime}),y)]$ (12)
B.2 Mixup
Mixup is the method that introduces interpolated samples to the training set Zhang et al. (2018b), which was proved to help model robustness and generalization Zhang & Deng (2021). For parameter $\lambda$ and samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$ randomly selected from the training set, the interpolated sample is:
$\widetilde{x}=\lambda x_{i}+(1-\lambda)x_{j},\quad\widetilde{y}=\lambda y_{i}+(1-\lambda)y_{j}$ (13)
Appendix C Adversarial Samples against PANN only
@@
1:  Initialize $\delta_{0}$ as all zero;
2:  while $\mathbb{F}(x+\delta)\neq y$ or $\widetilde{\mathbb{F}}(x+\delta)=y$ do
3:     if $\widetilde{\mathbb{F}}(x+\delta)=y$ then
4:        $\delta=\delta+\alpha\nabla_{\delta}\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)$
5:        $\delta=Clip_{[-\epsilon,\epsilon]}(Random\ Search(x+\delta))$
6:        $\delta=\delta\odot Mask(\nabla_{\delta}(\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta))\geq\epsilon_{atk})$
7:        $\delta=\delta\odot Mask(|\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x)|\leq\epsilon_{lim})$
8:     end if
9:     if $\mathbb{F}(x+\delta)\neq y$ then
10:        $Backtrack(\delta)$
11:     end if
12:  end while
13:  Return $\delta$

### v2

@@ Appendix A Additional Experiment Results
Appendix B Proof of Theorem 2.1 and 2.2
We first give two commonly used Theorem and Lemma
Theorem B.1.
[Mean Value Theorem for nondifferentiable functions.] Let $h:\mathbb{R}\to\mathbb{R}$ be a convex function and let $a<b$. Then there exists $c\in(a,b)$ such that $\frac{h(b)-h(a)}{b-a}\in\partial h(c)$.
Lemma B.2.
Let $h:\mathbb{R}\to\mathbb{R}$ be a convex function and let $\overline{z}\in\mathbb{R}$. Then $h$ has left derivative and right derivative at $\overline{z}$. Moreover,
$\mathop{\operatorname{sup}}_{z<\overline{z}}\phi_{\overline{z}}(z)=h^{{}^{\prime}}_{-}(\overline{z})\leq h^{{}^{\prime}}_{+}(\overline{z})=\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)$
where $\phi_{\overline{z}}(z)$ is defined as $\phi_{\overline{z}}(z)=\frac{h(z)-h(\overline{z})}{z-\overline{z}}$.
Lemma B.3 (Increased loss for $\overline{z}<0$).
Let function $h:\mathbb{R}\to\mathbb{R}$ be a convex function, and be differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\overline{z}\in\mathbb{R}$, $h^{\prime}_{+}(ReLU(\overline{z}))=0$ and $\overline{z}<0$, let $\epsilon\in\mathbb{R}$ be a small value. There exist an $\overline{\epsilon}\in(0,\epsilon)$ satisfies:
$0\leq\epsilon\cdot\mathop{\operatorname{sup}}_{z<\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)\leq\Delta h(\operatorname{ReLU}(\overline{z}))\leq\epsilon\cdot\mathop{\operatorname{inf}}_{z>\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)$ (25)
and
$\lim_{\epsilon\to 0}\Delta h(\operatorname{ReLU}(\overline{z}))=\epsilon\cdot\mathop{\operatorname{inf}}_{z>\overline{0}}\phi_{\overline{0}}(z)$ (26)
where $\Delta h(\operatorname{ReLU}(\overline{z}))=h(\operatorname{ReLU}(\overline{z})+\epsilon)-h(\operatorname{ReLU}(\overline{z}))$
Proof.
From $h^{\prime}(ReLU(\overline{z}))=0$ we can get $h(\operatorname{ReLU}(\overline{z})+\epsilon)\geq h(\operatorname{ReLU}(\overline{z}))$. From B.1 we know there exist a $\overline{\epsilon}\in(0,\epsilon)$ satisfies:
$h(\operatorname{ReLU}(\overline{z})+\epsilon)-h(\operatorname{ReLU}(\overline{z}))\in[\epsilon\cdot\mathop{\operatorname{sup}}_{z<\overline{\epsilon}}\phi_{\overline{\epsilon}}(z),\epsilon\cdot\mathop{\operatorname{inf}}_{z>\overline{\epsilon}}\phi_{\overline{0}}(z)]$ (27)
From Theorem B.1, it’s easy to get that for $z1<z2\in\mathbb{R}$, we have:
$\mathop{\operatorname{sup}}_{z<z_{1}}\phi_{z_{1}}(z)\leq\mathop{\operatorname{inf}}_{z>z_{1}}\phi_{z_{1}}(z)\leq\mathop{\operatorname{sup}}_{z<z_{2}}\phi_{z_{2}}(z)\leq\mathop{\operatorname{inf}}_{z>z_{2}}\phi_{z_{2}}(z)$ (28)
so we have:
$0\leq\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z)\leq\mathop{\operatorname{sup}}_{z<\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)\leq\mathop{\operatorname{inf}}_{z>\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)$ (29)
and:
$\lim_{\epsilon\to 0}\mathop{\operatorname{sup}}_{z<\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)=\lim_{\epsilon\to 0}\mathop{\operatorname{inf}}_{z>\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)=\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z)$ (30)
so we have:
$h(\operatorname{ReLU}(\overline{z})+\epsilon)-h(\operatorname{ReLU}(\overline{z}))\in[\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z),\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z)]$ (31)
which can get: $\Delta h(\operatorname{ReLU}(\overline{z}))=\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z)$
The proof is complete. ∎
Lemma B.4 (Increased loss for $\overline{z}>0$).
Let function $h:\mathbb{R}\to\mathbb{R}$ be a convex function, and be differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\overline{z}\in\mathbb{R}$, $h^{\prime}(ReLU(\overline{z}))=0$ and $\overline{z}>0$, let $\epsilon\in\mathbb{R}$ be a small value. We have:
$\lim_{\epsilon\to 0}\Delta h(\operatorname{ReLU}(\overline{z}))=0$ (32)
where $\Delta h(\operatorname{ReLU}(\overline{z}))=h(\operatorname{ReLU}(\overline{z})+\epsilon)-h(\operatorname{ReLU}(\overline{z}))$
Theorem B.5.
For two convex function $h_{1}:\mathbb{R}\to\mathbb{R}$ and $h_{2}:\mathbb{R}\to\mathbb{R}$ differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\overline{z}_{1}\in\mathbb{R}$ and $\overline{z}_{2}\in\mathbb{R}$ and $\overline{z}_{1}<0,\overline{z}_{2}>0$, such that $h_{1}^{\prime}(\overline{z}_{1})=h_{2}^{\prime}(\overline{z}_{2})=0$. Let $\epsilon\in\mathbb{R}$ be a small value, we have:
$\lim_{\epsilon\to 0}\Delta h_{1}-\Delta h_{2}=\epsilon\cdot\mathop{\operatorname{inf}}_{z>\overline{0},h_{1}}\phi_{\overline{0}}(z)$ (33)
Lemma B.6.
Let function $h:\mathbb{R}\to\mathbb{R}$ be a convex function, and be differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\overline{z}\in\mathbb{R}$ and $h^{\prime}(\operatorname{ReLU}(\overline{z}))=0$, let $\epsilon\in\mathbb{R}$ be a small value, we have:
$\epsilon\cdot\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)\leq\Delta h(\operatorname{ReLU}(\overline{z}))$ (34)
where $\Delta f(\operatorname{ReLU}(\overline{z}))=f(\operatorname{ReLU)(\overline{z}}+\epsilon)-f(\operatorname{ReLU}(\overline{z}))$
Proof.
For $\overline{z}>0$, $\mathop{\operatorname{sup}}_{z<\overline{z}}\phi_{\overline{z}}(z)=\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)=h^{\prime}(\overline{z})$, so we have:
$\epsilon\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)\leq\Delta h(\operatorname{ReLU}(\overline{z}))$ (35)
For $\overline{z}<0$, from Equation 29, we have:
$\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)\leq 0\leq\mathop{\operatorname{inf}}_{z>0}\phi_{0}(z)\leq\mathop{\operatorname{sup}}_{z<\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)\leq\mathop{\operatorname{inf}}_{z>\overline{\epsilon}}\phi_{\overline{\epsilon}}(z)$ (36)
so we can also get:
$\epsilon\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)\leq\Delta h(\operatorname{ReLU}(\overline{z}))$ (37)
The proof is complete. ∎
And from [12] we have the theorem:
Theorem B.7.
For an $\mathcal{L}$-smooth function $L(W)$, assume $L$ is lower-bounded as $L(W)\geq L^{\star}$, $L(W,X)$ is the loss over one minibatch $X$, $\sigma^{2}$ is the gradient noise variance, $\mathbb{E}[\nabla L(W,X)-\nabla L(W)]=0$, $\mathbb{E}[\|\nabla L(W,X)-\nabla L(W)\|^{2}]\leq\delta^{2}$, and $\|\nabla L(W)\|\leq G$ for any $W$. If $\eta\leq\frac{C}{\sqrt{t+1}}$:
$\min_{k=0,\ldots,t}\mathbb{E}[\|\nabla f(W_{k})\|^{2}]\leq\frac{1}{\sqrt{t+1}}\left[C_{1}+C_{2}\right],$ (38)
where
$C_{1}$ $=\frac{L(W_{0})+\frac{\lambda}{2}\|W_{0}\|^{2}-L^{\star}}{C},$ (39)
$C_{2}$ $=C(\mathcal{L}+\lambda)((G+\sup(\lambda\|W\|))^{2}+\sigma^{2}),$ (40)
Theorem B.8.
Let function $L:\mathbb{R}\to\mathbb{R}$ be a $\mathcal{L}$-smooth function and $h$ be a convex function with $h^{\prime}(\operatorname{ReLU}(\overline{z}))=0$. Assume $h$ is differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\epsilon\in\mathbb{R}$ be a small value. $\mathbb{E}[\nabla L(\textbf{{W}},X)-\nabla L(\textbf{{W}})]=0$, $\mathbb{E}[\|\nabla L(\textbf{{W}},X)-\nabla L(\textbf{{W}})\|^{2}]\leq\sigma^{2}$, $\|\nabla l(\textbf{{W}})\|\leq G$ for any W, and $\eta\leq\frac{C}{\sqrt{t+1}}$. If the model is trained $t^{\prime}+1$ epochs after it has reached the stationary point $\textbf{{W}}^{*}$ where $L(\textbf{{W}}^{*},X)=L^{\star}$, the increased loss $\Delta L=h(\operatorname{ReLU}(\overline{z}+\epsilon))-h(\operatorname{ReLU}(\overline{z}))$ caused by error $\epsilon$ is bounded by:
$||\epsilon||^{2}\cdot\mathbb{E}[||\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]\leq\mathbb{E}[||\Delta h||^{2}]$ (41)
where
$\operatorname{sup}\mathbb{E}[||\operatorname{inf}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]\leq||h^{\prime}_{+,t}(\overline{z})||^{2}+\frac{t^{\prime}(C_{1}+C_{2})}{d\sqrt{t+t^{\prime}+1}}$ (42)
and:
$C_{1}$ $=\frac{L(\textbf{{W}}_{0})-L^{\star}}{C}$ (43)
$C_{2}$ $=C(\mathcal{L}+\lambda)((G+\sup(\lambda\|W\|))^{2}+\sigma^{2}),$ (44)
Proof.
From lemma B.6, we have:
$\epsilon\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)\leq\Delta h(\operatorname{ReLU}(\overline{z}))$ (45)
It’s easy to get
$\operatorname{sup}\mathbb{E}[||\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]=||h^{\prime}_{+,t}(\overline{z})||^{2}+\frac{1}{d}\mathbb{E}[\sum_{i=t}^{t+t^{\prime}}||\nabla f(\operatorname{ReLU}(z))||^{2}]$ (46)
From theorem B.7 we have:
$\sup\mathbb{E}[\sum\limits_{i=t}^{t+t^{\prime}+1}||\nabla f(\operatorname{ReLU}(z))||^{2}]=\mathop{\operatorname{\min}}_{k=0,\ldots,t+t^{\prime}+1}(t^{\prime}+1)\mathbb{E}[\|\nabla f(W_{k})x\|^{2}]\leq\frac{(t^{\prime}+1)C_{0}}{\sqrt{t+t^{\prime}+1}}$ (47)
such that we have:
$\operatorname{sup}\mathbb{E}[||\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]=||h^{\prime}_{+,t}(\overline{z})||^{2}+\frac{1}{d}\sup\mathbb{E}[\sum\limits_{i=t}^{t+t^{\prime}+1}||\nabla f(\operatorname{ReLU}(z))||^{2}]\leq||f^{\prime}_{+,t}(\overline{z})||^{2}+\frac{(t^{\prime}+1)C_{0}}{d\sqrt{t+t^{\prime}+1}}$ (48)
The proof is complete. ∎
Appendix C Proof of theorem B.7
Here we borrow the proof from [12] for reader’s reference:
Lemma C.1 (Convergence of SGD, a specialized case of Theorem 1 in [41] with $\beta=s=0$).
Assume that $L(\theta)$ is an $\mathcal{L}$-smooth function [^3] It means that $\|\nabla L(\theta_{a})-\nabla L(\theta_{b})\|\leq\mathcal{L}\|\theta_{a}-\theta_{b}\|$ holds for any $\theta_{a}$ and $\theta_{b}$., $L$ is lower bounded as $L(\theta)\geq L^{\star}$, $\mathbb{E}[\nabla L(\theta,X)-\nabla L(\theta)]=0$, $\mathbb{E}[\|\nabla L(\theta,X)-\nabla L(\theta)\|^{2}]\leq\delta^{2}$, $\|\nabla L(\theta)\|\leq G$ for any $\theta$. Let SGD optimize $L$ for $t+1$ iterations. If $\eta\leq\frac{C}{\sqrt{t+1}}$, we have
$\min_{k=0,\ldots,t}\mathbb{E}[\|\nabla L(\theta_{k})\|^{2}]\leq\frac{C_{0}}{\sqrt{t+1}},$ (49)
where $C_{0}=\left[\frac{L(\theta_{0})-L^{\star}}{C}+C\mathcal{L}(G^{2}+\sigma^{2})\right]$.
Proof.
Given the conditions of $L(\theta)$ in Lemma C.1, we may obtain the resulted conditions of $f(\theta)=L(\theta)+\frac{\lambda}{2}\|\theta\|^{2}$.
As $L(\theta)$ is an $\mathcal{L}$-smooth function, we have
$\|\nabla f(\theta_{a})-\nabla f(\theta_{b})\|=\|\nabla L(\theta_{a})-\nabla L(\theta_{b})+\lambda(\theta_{a}-\theta_{b})\|\leq(\mathcal{L}+\lambda)\|\theta_{a}-\theta_{b}\|$ (50)
holds for any $\theta_{a}$ and $\theta_{b}$. It shows that $f(\theta)$ is an $(\mathcal{L}+\lambda)$-smooth function.
As $L$ is lower bounded as $L(\theta)\geq L^{\star}$, we have
$f^{\star}\geq L^{\star}.$ (51)
As $\mathbb{E}[\nabla L(\theta,X)-\nabla L(\theta)]=0$, we have
$\mathbb{E}[\nabla f(\theta,X)-\nabla f(\theta)]=\mathbb{E}[\nabla L(\theta,X)-\nabla L(\theta)]=0.$ (52)
As $\mathbb{E}[\|\nabla L(\theta,X)-\nabla L(\theta)\|^{2}]\leq\delta^{2}$, we have
$\mathbb{E}[\|\nabla f(\theta,X)-\nabla f(\theta)\|^{2}]=\mathbb{E}[\|\nabla L(\theta,X)-\nabla L(\theta)\|^{2}]\leq\delta^{2}.$ (53)
As $\|\nabla L(\theta)\|\leq G$, we have
$\|\nabla f(\theta)\|=\|\nabla L(\theta)+\lambda\theta\|\leq G+\lambda\|\theta\|_{\max},$ (54)
where $\|\theta\|_{\max}$ is the maximum $L_{2}$ norm of any $\theta$.
Introducing the derived conditions Eq. (12) - (16) for $f$ into Lemma C.1, we may treat $f$ as the objective optimized by SGD. Then we have
$\min_{k=0,\ldots,t}\mathbb{E}[\|\nabla f(\theta_{k})\|^{2}]$ $\leq\frac{1}{\sqrt{t+1}}\left[\frac{f(\theta_{0})-f^{\star}}{C}+C(\mathcal{L}+\lambda)((G+\lambda\|\theta\|_{\max})^{2}+\sigma^{2})\right]$ (55)
$\leq\frac{1}{\sqrt{t+1}}\left[\frac{L(\theta_{0})+\frac{\lambda}{2}\|\theta_{0}\|^{2}-L^{\star}}{C}+C(\mathcal{L}+\lambda)((G+\lambda\|\theta\|_{\max})^{2}+\sigma^{2})\right]$ (56)
Obviously, the gradient norm upper bound in convergence analysis monotonically increases as the weight decay strength $\lambda$.
The proof is complete.
Appendix D Adversarial Samples against PANN only
@@
  Initialize $\delta_{0}$ as all zero;
  while $\mathbb{F}(x+\delta)\neq y$ or $\widetilde{\mathbb{F}}(x+\delta)=y$ do
     if $\widetilde{\mathbb{F}}(x+\delta)=y$ then
        $\delta=\delta+\alpha\nabla_{\delta}\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)$
        $\delta=Clip_{[-\epsilon,\epsilon]}(Random\ Search(x+\delta))$
        $\delta=\delta\odot Mask(\nabla_{\delta}(\mathcal{L}_{\widetilde{\mathbb{F}}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta))\geq\epsilon_{atk})$
        $\delta=\delta\odot Mask(|\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x+\delta)-\nabla_{\delta}\mathcal{L}_{\mathbb{F}}(x)|\leq\epsilon_{lim})$
     end if
     if $\mathbb{F}(x+\delta)\neq y$ then
        $Backtrack(\delta)$
     end if
  end while
  Return $\delta$